OpenStudy (anonymous):

What is the slope of a line that with the points (2,4) and (4,-2)? 1 -1/3 3 -3

OpenStudy (anonymous):

Slope=y2−y1x2−x1 lolol the guy gave you the equation in your last post

OpenStudy (anonymous):

(y2-y1) / (x2-x1)**

OpenStudy (anonymous):

slope = (-2-4)/(4-2) = -6/2=-3